Rudd Clothiers is a small company that manufactures tall-men’s suits. The company has used a standard cost accounting system. In May 2020, 10,100 suits were produced. The following standard and actual cost data applied to the month of May when normal capacity was 14,000 direct labor hours. All materials purchased were used.

Cost Element

Standard (per unit)

Actual

Direct materials 7 yards at $4.20 per yard $292,330 for 71,300 yards ($4.10 per yard)
Direct labor 1.10 hours at $14.00 per hour $167,453 for 11,710 hours ($14.30 per hour)
Overhead 1.10 hours at $6.30 per hour (fixed $3.50; variable $2.80) $48,900 fixed overhead $37,500 variable overhead


Overhead is applied on the basis of direct labor hours. At normal capacity, budgeted fixed overhead costs were $49,000, and budgeted variable overhead was $39,200.

(a)

Compute the total, price, and quantity variances for (1) materials and (2) labor. (Round per unit values to 2 decimal places, e.g. 52.75 and final answers to 0 decimal places, e.g. 52.)

(1) Total materials variance $
Materials price variance $
Materials quantity variance $
(2) Total labor variance $
Labor price variance $
Labor quantity variance $


(b)

Compute the total overhead variance.

Total overhead variance $
Answer #1

1. Total material variance = standard cost - actual cost

= 10100x7x4.20 -292330 = $4610 Favorable

Materials price variance = (Standard price - actual price) x Material used
= {$4.2 - 4.10) x 71300 = $7130 Favorable

Materials quantity variance = (Standard quantity - actual quantity used) x standard price
= {(10100x7) - 71300} x $4.2 = $2520 Unfavorabler

2 total labor cost variance = 10100x1.1x14 - 167453 = 11913 unfavorable

. labor rate variance = (Standard rate - actual rate) x Actual hours used
= ($14 - 14.3) x 11710 = $3513 Unfavorable

Labor efficiency variance = (Standard hours - actual hours) x Standard rate
= {(10100x1.1) - 11710} x $14.0= $8400 Unfavorable

b. Total overhead variance = standard cost - actual cost

= 10100x1.1x6.30 - 48900+37500

= $16407 Unfavorable